# Provable Sample-Efficient Transfer Learning Conditional Diffusion Models via Representation Learning

### v1

@@ 5.2 Behavior Cloning via Meta-Diffusion Policy
$V_{M}(\pi,s_{0}):=\mathbb{E}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r_{M}(s_{t},a_{t})\Big{]},a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim\mathcal{T}_{M}(\cdot|s_{t},a_{t}),$ (5.4)
$V_{M}(\pi):=\mathbb{E}_{s_{0}\sim\rho}[V_{M}(\pi,s_{0})].$
@@
[…]
@@
Theorem 5.2.
Suppose the expert policy $\pi_{*}^{k}$ satisfies Assumption 3.1, 3.2, 3.3. Then under meta-learning setting, it holds that with probability no less than $1-\delta$,
$\mathbb{E}_{M^{0}}\mathbb{E}_{\{(s_{i}^{0},a_{i}^{0})\}_{i=1}^{m}\sim d_{*}^{0}}[V_{M^{0}}(\pi_{*}^{0})-V_{M^{0}}(\widehat{\pi}^{0})]\lesssim\frac{1}{(1-\gamma)^{2}}\left[\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}(m\wedge n)}{(m\wedge n)^{\frac{1}{d_{a}+d_{s}+9}}}+\frac{\log^{2}(nK/\delta)}{K^{\frac{1}{D_{s}+2}}}\right].$ (5.5)
@@
$\mathbb{E}_{\{(s_{i}^{0},a_{i}^{0})\}_{i=1}^{m}\sim d_{*}^{0}}[V_{M^{0}}(\pi_{*}^{0})-V_{M^{0}}(\widehat{\pi}^{0})]\lesssim\frac{1}{(1-\gamma)^{2}}\left[\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}((m/\nu)\wedge n)}{\nu^{\frac{1}{2}}((m/\nu)\wedge n)^{\frac{1}{d_{a}+d_{s}+9}}}+\frac{\log^{2}(nK/\delta)}{\nu^{\frac{1}{2}}(nK)^{\frac{1}{D_{s}+2}}}+\sqrt{\Delta}\right].$ (5.6)
The proof can be found in Appendix C.2. This provides the first statistical guarantee of diffusion policy in behavior cloning. Notably, in both cases, the number of source tasks $K$ has an exponential dependence on $D_{s}$, further suggesting the importance of data coverage when tackling distribution shift in offline meta-RL [Pong et al., 2022].
6 Experiments
Experimental Settings
Our numerical example is the high-dimensional conditioned diffusion [Cui et al., 2016; Yu et al., 2023] arising from the following Langevin SDE
where $\beta>0$ and $w_{s}$ is a one-dimensional standard Brownian motion. The SDE (6.1) is discretized by the Euler-Maruyama scheme with a step size of 0.020.02 0.020.02, which defines the prior distribution $p_{\beta}(x)$ for the (discretized) trajectory $x=\left(u_{0.02},u_{0.04},\ldots,u_{1.00}\right)^{\top}\in\mathbb{R}^{50}$. We consider a conditional Gaussian likelihood function, $p(y|x)=\mathcal{N}(Mx,I_{100}/4)$, where $M\in\mathbb{R}^{100\times 50}$ is a pre-defined projection matrix. With a set of pre-selected $\{\beta_{k};1\leq k\leq K\}$ with $\beta_{k}=k$ and $K=10$, the $k$-th source distribution of $(x,y)$ is given by $\mathbb{P}^{k}(x,y)=p_{\beta_{k}}(x)p(y|x)$. The target distribution $\mathbb{P}^{0}(x,y)$ is given by $\beta_{0}=5.5$ or $\beta_{0}=15$.
Each $f^{k}$ and $f^{0}$ are implemented as a 2-layer MLP with 128 internal channels and 60 input channels. The representation map $h$ is implemented as a 5-layer MLP with 512 internal channels and 10 output channels. We have $n=1000$ pre-training samples from each source distribution $\mathbb{P}^{k}$, $m\in\{10,20,30,40,50,100\}$ fine-tuning samples from the target distribution $\mathbb{P}^{0}$, and 100100 100100 test samples from the target distribution $\mathbb{P}^{0}$ for evaluating the test error of different models. In the pre-training phase, the ${\widehat{f}^{k};1\leq k\leq K}$ and $\hat{h}$ are trained on the $K=10$ source distributions with 400K iterations and a batch size of 512. In the fine-tuning phase, the pre-trained representation map $\widehat{h}$ is fixed, and the $\widehat{f}^{0}$ is trained on the target distribution with 200K iterations and a batch size of $m$. As an important baseline, we also consider jointly training $h$ and $f^{0}$ on the target distribution from scratch, using the same fine-tuning samples.
Results
We report the MSEs of the estimated posterior mean of $\mathbb{P}^{0}(x|y)$ on the test samples in Table 2 and 3. We see that for different $\beta$’s and $m$’s, the fine-tuned models can provide significantly more accurate posterior mean estimations in almost all of the cases, suggesting the effectiveness of the representation map $\widehat{h}$ learned in the pre-training phase. We also notice a large variance among the results of different replicates, and attribute the slightly worse performance of fine-tined models at $m=50,\beta=5.5$ to the potential randomness. As $m$ increases, the performance gaps between fine-tuned models and train-from-scratch models get smaller, since more training samples yield more generalization benefits.
$m$ 10 20 30 40 50 100
MSE of fine-tuned models 14.47 3.68 2.45 1.82 1.9 0.91
MSE of train-from-scratch models 21.99 10.61 5.71 2.38 1.77 1.04
Table 2: The MSEs of different models for estimating the posterior mean of $x$, whose ground truth is formed by an extremely long LMC run ($\beta_{0}=5.5$).
$m$ 10 20 30 40 50 100
MSE of fine-tuned models 6.14 2.65 1.61 1.08 0.96 0.45
MSE of train-from-scratch models 24.41 20.62 18.67 13.49 7.03 1.23
Table 3: The MSEs of different models for estimating the posterior mean of $x$, whose ground truth is formed by an extremely long LMC run ($\beta_{0}=15$).
7 Conclusion
In this paper, we take the first step towards understanding the sample efficiency of transfer learning conditional diffusion models from the perspective of representation learning. We provide generalization guarantee of transferring score matching in CDMs in different settings. We further establish an end-to-end distribution estimation error bound using deep neural network. Two practical applications are investigated based on our theoretical results. We hope this work can motivate future theoretical study on the popular transfer learning paradigm in generative AIs.
Appendix A Proofs in Section 3
A.1 Preliminaries
Lemma A.1.
@@ A.1 Preliminaries
$p_{t}(x|y)=\int\phi_{t}(x|x_{0})p(x_{0}|y)\mathrm{d}x_{0},\quad\phi_{t}(x|x_{0})=\frac{1}{(2\pi\sigma_{t}^{2})^{\frac{d_{x}}{2}}}\exp\Big{(}-\frac{\|x-\alpha_{t}x_{0}\|^{2}}{2\sigma_{t}^{2}}\Big{)}.$ (A.1)
@@
$\displaystyle\nabla_{x}\log p_{t}(x|y)$ $\displaystyle=\int\nabla_{x}\log\phi_{t}(x|x_{0})\frac{\phi_{t}(x|x_{0})p(x_{0}|y)}{\int\phi_{t}(x|z)p(z|y)\mathrm{d}z}\mathrm{d}x_{0}$ (A.2)
$\displaystyle=\frac{1}{\alpha_{t}}\int\nabla_{x}\log p(x_{0}|y)\frac{\phi_{t}(x|x_{0})p(x_{0}|y)}{\int\phi_{t}(x|z)p(z|y)\mathrm{d}z}\mathrm{d}x_{0}.$ (A.3)
@@
(A.1) can be directly implied by the definition of forward process. And it yields
$\nabla_{x}\log p_{t}(x|y)$ $=\frac{\nabla_{x}p_{t}(x|y)}{p_{t}(x|y)}$ (A.4)
@@
which is (A.2). Moreover, noticing that $\nabla_{x}\phi_{t}(x|x_{0})=-\frac{1}{\alpha_{t}}\nabla_{x_{0}}\phi_{t}(x|x_{0})$, then by integration by parts,
$\frac{\int\nabla_{x}\phi_{t}(x|x_{0})p(x_{0}|y)\mathrm{d}x_{0}}{\int\phi_{t}(x|x_{0})p(x_{0}|y)\mathrm{d}x_{0}}$ $=-\frac{1}{\alpha_{t}}\frac{\int\nabla_{x_{0}}\phi_{t}(x|x_{0})p(x_{0}|y)\mathrm{d}x_{0}}{\int\phi_{t}(x|x_{0})p(x_{0}|y)\mathrm{d}x_{0}}$ (A.5)
@@
Hence (A.3) is proved. ∎
Lemma A.2.
[Lem. 3.1] For any $w\in[0,1]^{d_{y}}$, denote the conditional score of forward process $\nabla_{x}\log p_{t}(x;w)$ by $f_{*}(x,w,t)$. Then there exist constants $C_{X},C_{X}^{\prime}$, such that for any $t\in[0,T]$, the function $f_{*}(x,w,t)$ is $(C_{X}+C_{X}^{\prime}\|x\|^{2})$-Lipschitz in $x$, $(C_{X}+C_{X}^{\prime}\|x\|)$-Lipschitz in $w$.
@@
When $t\geq\frac{1}{2(L+1)}$, according to (A.2), we have
$\nabla_{x}f_{*}(x,w,t)$ $=\nabla_{x}^{2}\log p_{t}(x;w)$ (A.6)
@@
$\mathrm{Var}_{q_{t}(x_{0}|x,w)}\Big{(}\frac{\alpha_{t}x_{0}-x}{\sigma_{t}^{2}}\Big{)}$ $\preceq\frac{1}{\sigma_{t}^{2}}\int\big{\|}\frac{\alpha_{t}x_{0}-x}{\sigma_{t}}\big{\|}^{2}\frac{\phi_{t}(x|x_{0})p(x_{0}|y)}{\int\phi_{t}(x|z)p(z|y)\mathrm{d}z}\mathrm{d}x_{0}$ (A.7)
@@
$\int_{\|\frac{\alpha_{t}x_{0}-x}{\sigma_{t}}\|\leq R/2}p(x_{0};w)\mathrm{d}x_{0}\geq\int_{\|x_{0}\|\leq C_{0}}p(x_{0};w)\mathrm{d}x_{0}\geq 1-2\exp(-C_{1}^{\prime}C_{0}^{2})\geq\frac{1}{2},$ (A.8)
@@
$\|\nabla_{x}f_{*}(x,w,t)\|\leq\frac{1}{\sigma_{t}^{2}}+\big{\|}\mathrm{Var}_{q_{t}(x_{0}|x,w)}\Big{(}\frac{\alpha_{t}x_{0}-x}{\sigma_{t}^{2}}\Big{)}\big{\|}\leq\frac{R^{2}}{\sigma_{t}^{2}}+\frac{2}{\sigma_{t}^{2}}\leq\frac{8\|x\|^{2}+8C_{0}^{2}+2\sigma_{t}^{2}}{\sigma_{t}^{4}}.$ (A.9)
@@
$\nabla_{w}f_{*}(x,w,t)$ $=\mathrm{Cov}_{q_{t}(x_{0}|x,w)}\big{(}\nabla_{x}\log\phi_{t}(x|x_{0}),\nabla_{w}\log p(x_{0};w)\big{)}$ (A.10)
@@
$\|\nabla_{w}f_{*}(x,w,t)\|$ $\leq B\sqrt{\big{\|}\mathrm{Var}_{q_{t}(x_{0}|x,w)}\Big{(}\frac{\alpha_{t}x_{0}-x}{\sigma_{t}^{2}}\Big{)}\big{\|}}$ (A.11)
@@
=(A.3)${=}\frac{1}{\alpha_{t}^{2}}\mathbb{E}_{q_{t}(x_{0}|x,y)}\left[\nabla_{x}^{2}\log p(x_{0};w)\right]+\frac{1}{\alpha_{t}^{2}}\mathrm{Var}_{q_{t}(x_{0}|x,w)}\big{(}\nabla_{x}\log p(x_{0};w)\big{)}.$ ) .
@@
$\mathrm{Var}_{q_{t}(x_{0}|x,w)}\big{(}\nabla_{x}\log p(x_{0};w)\big{)}\preceq L^{-1}\mathbb{E}\left[\nabla_{x}^{2}\log p(x_{0};w)(\nabla_{x}^{2}\log p(x_{0};w))^{\top}\right]\leq L.$ (A.13)
@@
$\|\nabla_{x}f_{*}(x,w,t)\|\leq\frac{2L}{\alpha_{t}^{2}}.$ (A.14)
@@
$\nabla_{w}f_{*}(x,w,t)$ $=\frac{1}{\alpha_{t}}\mathbb{E}_{q_{t}(x_{0}|x,w)}\left[\nabla_{w}\nabla_{x}\log p(x_{0};w)\right]+\frac{1}{\alpha_{t}}\mathrm{Cov}_{q_{t}(x_{0}|x,w)}\big{(}\nabla_{x}\log p(x_{0};w),\nabla_{w}\log p(x_{0};w)\big{)}$ (A.15)
@@
Combine all the arguments in (A.9),(A.11),(A.14),(A.15) and we complete the proof. ∎
Lemma A.3 (Lemma 7, Chen et al. [2022a]).
@@
$\log\mathcal{N}(\mathcal{F},\|\cdot\|_{L^{\infty}([-R,R]^{d_{x}+d_{y}+1})},\varepsilon)\lesssim S_{f}L_{f}\log\left(\frac{L_{f}W_{f}(B_{f}\vee 1)RM_{f}}{\varepsilon}\right).$ (A.16)
@@
$\log\mathcal{N}(\mathcal{H},\|\cdot\|_{L^{\infty}([0,1]^{D_{y}})},\varepsilon)\lesssim S_{h}L_{h}\log\left(\frac{L_{h}W_{h}(B_{h}\vee 1)}{\varepsilon}\right).$ (A.17)
A.2 Proofs of Transfer Learning
Proposition A.4 (Prop. 3.2).
[…]
$\widehat{f}:=\operatorname*{arg\,min}_{f\in\mathcal{F}}\frac{1}{m}\sum_{i=1}^{m}\ell(x_{i},y_{i},s_{f,h}).$ (A.18)

### v2

@@ A.2 Behavior Cloning via Meta-Diffusion Policy
$V_{M}(\pi,s_{0})=\mathbb{E}\Big[\sum_{t=0}^{\infty}\gamma^{t}r_{M}(s_{t},a_{t})\Big],a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim\mathcal{T}_{M}(\cdot|s_{t},a_{t}),$ (A.4)
$V_{M}(\pi)=\mathbb{E}_{s_{0}\sim\rho}[V_{M}(\pi,s_{0})].$
@@
Suppose there are $K$ source tasks $M^{1},\cdots,M^{K}\in\mathcal{M}$, and the expert policy of each task is denoted as $\pi_{*}^{k}$. In behavior cloning, for each source task $M^{k}$, we have $n$ pairs of $\{(s_{i}^{k},a_{i}^{k})\}_{i=1}^{n}\overset{\textit{i.i.d.}}{\sim}d^{k}_{*}:=d_{M^{k}}^{\pi_{*}^{k}}$. The goal is to imitate the expert policy of target task $M^{0}\in\mathcal{M}$, of which the sample size is only $m\ll n$.
@@
Theorem A.2.
Suppose the expert policy $\pi_{*}^{k}$ satisfies Assumption 3.1, 3.2, $3.$ Then under meta-learning setting, it holds that with probability no less than $1-\delta$,
$\mathbb{E}_{M^{0}}\mathbb{E}_{\{(s_{i}^{0},a_{i}^{0})\}_{i=1}^{m}\sim d_{*}^{0}}[V_{M^{0}}(\pi_{*}^{0})-V_{M^{0}}(\widehat{\pi}^{0})]\lesssim\frac{1}{(1-\gamma)^{2}}\left[\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}(m\wedge n)}{(m\wedge n)^{\frac{1}{d_{a}+d_{s}+9}}}+\frac{\log^{2}(nK/\delta)}{K^{\frac{1}{D_{s}+2}}}\right].$ (A.5)
@@
$\mathbb{E}_{\{(s_{i}^{0},a_{i}^{0})\}_{i=1}^{m}\sim d_{*}^{0}}[V_{M^{0}}(\pi_{*}^{0})-V_{M^{0}}(\widehat{\pi}^{0})]\lesssim\frac{1}{(1-\gamma)^{2}}\left[\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}((m/\nu)\wedge n)}{\nu^{\frac{1}{2}}((m/\nu)\wedge n)^{\frac{1}{d_{a}+d_{s}+9}}}+\frac{\log^{2}(nK/\delta)}{\nu^{\frac{1}{2}}(nK)^{\frac{1}{D_{s}+2}}}+\sqrt{\Delta}\right].$ (A.6)
The proof can be found in Appendix D.2. This provides the first statistical guarantee of diffusion policy in behavior cloning. Notably, in both cases, the number of source tasks $K$ has an exponential dependence on $D_{s}$, further suggesting the importance of data coverage when tackling distribution shift in offline meta-RL [Pong et al., 2022].
Appendix B Proofs in Section 3
B.1 Preliminaries
Lemma B.1.
@@ B.1 Preliminaries
$p_{t}(x|y)=\int\phi_{t}(x|x_{0})p(x_{0}|y)\mathrm{d}x_{0},\quad\phi_{t}(x|x_{0})=\frac{1}{(2\pi\sigma_{t}^{2})^{\frac{d_{x}}{2}}}\exp\Big(-\frac{\|x-\alpha_{t}x_{0}\|^{2}}{2\sigma_{t}^{2}}\Big).$ (B.1)
@@
$\displaystyle\nabla_{x}\log p_{t}(x|y)$ $\displaystyle=\int\nabla_{x}\log\phi_{t}(x|x_{0})\frac{\phi_{t}(x|x_{0})p(x_{0}|y)}{\int\phi_{t}(x|z)p(z|y)\mathrm{d}z}\mathrm{d}x_{0}$ (B.2)
$\displaystyle=\frac{1}{\alpha_{t}}\int\nabla_{x}\log p(x_{0}|y)\frac{\phi_{t}(x|x_{0})p(x_{0}|y)}{\int\phi_{t}(x|z)p(z|y)\mathrm{d}z}\mathrm{d}x_{0}.$ (B.3)
@@
(B.1) can be directly implied by the definition of forward process. And it yields
$\nabla_{x}\log p_{t}(x|y)$ $=\frac{\nabla_{x}p_{t}(x|y)}{p_{t}(x|y)}$ (B.4)
@@
which is (B.2). Moreover, noticing that $\nabla_{x}\phi_{t}(x|x_{0})=-\frac{1}{\alpha_{t}}\nabla_{x_{0}}\phi_{t}(x|x_{0})$, then by integration by parts,
$\frac{\int\nabla_{x}\phi_{t}(x|x_{0})p(x_{0}|y)\mathrm{d}x_{0}}{\int\phi_{t}(x|x_{0})p(x_{0}|y)\mathrm{d}x_{0}}$ $=-\frac{1}{\alpha_{t}}\frac{\int\nabla_{x_{0}}\phi_{t}(x|x_{0})p(x_{0}|y)\mathrm{d}x_{0}}{\int\phi_{t}(x|x_{0})p(x_{0}|y)\mathrm{d}x_{0}}$ (B.5)
@@
Hence (B.3) is proved. ∎
Lemma B.2.
[Lem. 3.1] For any $w\in[0,1]^{d_{y}}$, denote the conditional score of forward process $\nabla_{x}\log p_{t}(x;w)$ by $f_{*}(x,w,t)$. Then there exist constants $C_{X},C_{X}^{\prime}$, such that for any $R>0$, the function $f_{*}(x,w,t)$ is $(C_{X}+C_{X}^{\prime}R^{2})$-Lipschitz in $x$, $(C_{X}+C_{X}^{\prime}R)$-Lipschitz in $w$, in the domain $\mathcal{B}_{R}\times[0,1]^{d_{y}}\times[0,T]$. Here $\mathcal{B}_{R}$ denotes the ball with radius $R$ centered at the origin.
@@
When $t\geq\frac{1}{2(L+1)}$, according to (B.2), we have
$\nabla_{x}f_{*}(x,w,t)$ $=\nabla_{x}^{2}\log p_{t}(x;w)$ (B.6)
@@
$\mathrm{Var}_{q_{t}(x_{0}|x,w)}\Big(\frac{\alpha_{t}x_{0}-x}{\sigma_{t}^{2}}\Big)$ $\preceq\frac{1}{\sigma_{t}^{2}}\int\big\|\frac{\alpha_{t}x_{0}-x}{\sigma_{t}}\big\|^{2}\frac{\phi_{t}(x|x_{0})p(x_{0}|y)}{\int\phi_{t}(x|z)p(z|y)\mathrm{d}z}\mathrm{d}x_{0}$ (B.7)
@@
$\|\nabla_{x}f_{*}(x,w,t)\|\leq\frac{1}{\sigma_{t}^{2}}+\big\|\mathrm{Var}_{q_{t}(x_{0}|x,w)}\Big(\frac{\alpha_{t}x_{0}-x}{\sigma_{t}^{2}}\Big)\big\|\leq\frac{R^{2}}{\sigma_{t}^{2}}+\frac{2}{\sigma_{t}^{2}}\leq\frac{8\|x\|^{2}+8C_{0}^{2}+2\sigma_{t}^{2}}{\sigma_{t}^{4}}.$ (B.9)
@@
$\nabla_{w}f_{*}(x,w,t)$ $=\mathrm{Cov}_{q_{t}(x_{0}|x,w)}\big(\nabla_{x}\log\phi_{t}(x|x_{0}),\nabla_{w}\log p(x_{0};w)\big)$ (B.10)
@@
$\|\nabla_{w}f_{*}(x,w,t)\|$ $\leq B\sqrt{\big\|\mathrm{Var}_{q_{t}(x_{0}|x,w)}\Big(\frac{\alpha_{t}x_{0}-x}{\sigma_{t}^{2}}\Big)\big\|}$ (B.11)
@@
[…]
@@
$\mathrm{Var}_{q_{t}(x_{0}|x,w)}\big(\nabla_{x}\log p(x_{0};w)\big)\preceq L^{-1}\mathbb{E}\left[\nabla_{x}^{2}\log p(x_{0};w)(\nabla_{x}^{2}\log p(x_{0};w))^{\top}\right]\leq L.$ (B.13)
@@
$\|\nabla_{x}f_{*}(x,w,t)\|\leq\frac{2L}{\alpha_{t}^{2}}.$ (B.14)
@@
$\nabla_{w}f_{*}(x,w,t)$ $=\frac{1}{\alpha_{t}}\mathbb{E}_{q_{t}(x_{0}|x,w)}\left[\nabla_{w}\nabla_{x}\log p(x_{0};w)\right]+\frac{1}{\alpha_{t}}\mathrm{Cov}_{q_{t}(x_{0}|x,w)}\big(\nabla_{x}\log p(x_{0};w),\nabla_{w}\log p(x_{0};w)\big)$ (B.15)
@@
Combine all the arguments in (B.9),(B.11),(B.14),(B.15) and we complete the proof. ∎
Lemma B.3 (Lemma 7, Chen et al. [2022a]).
@@
$\log\mathcal{N}(\mathcal{F},\|\cdot\|_{L^{\infty}([-R,R]^{d_{x}+d_{y}+1})},\varepsilon)\lesssim S_{f}L_{f}\log\left(\frac{L_{f}W_{f}(B_{f}\vee 1)R}{\varepsilon}\right).$ (B.16)
@@
$\log\mathcal{N}(\mathcal{H},\|\cdot\|_{L^{\infty}([0,1]^{D_{y}})},\varepsilon)\lesssim S_{h}L_{h}\log\left(\frac{L_{h}W_{h}(B_{h}\vee 1)}{\varepsilon}\right).$ (B.17)
B.2 Proofs of Transfer Learning
Proposition B.4 (Prop. 3.2).
Under Assumption 3.1, 3.2, 3.3, there exists some constant $C_{xy}$ such that the following holds. For any $h\in\mathcal{H}$ and $(x_{1},y_{1}),\cdots,(x_{m},y_{m})\overset{\textit{i.i.d.}}{\sim}\mathbb{P}$, define the empirical minimizer
$\widehat{f}:=\operatorname*{arg\,min}_{f\in\mathcal{F}}\frac{1}{m}\sum_{i=1}^{m}\ell(x_{i},y_{i},s_{f,h}).$ (B.18)